# Scaling-Up the Pretraining of the Earth Observation Foundation Model PhilEO to the MajorTOM Dataset

### v1

@@ Abstract
Today, Earth Observation (EO) satellites generate massive volumes of data, with the Copernicus Sentinel-2 constellation alone producing approximately 1.6TB per day. To fully exploit this information, it is essential to pretrain EO Foundation Models (FMs) on large unlabeled datasets, enabling efficient fine-tuning for downstream tasks with minimal labeled data. In this work, we present the scaling-up of our recently proposed EO Foundation Model, PhilEO Geo-Aware U-Net, on the unlabeled 23TB dataset MajorTOM, which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. We develop and study various PhilEO models with different numbers of parameters and architectures. Finally, we fine-tune the models on the PhilEO Bench for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance. Our results demonstrate that for all n-shots for road density regression, the PhilEO 44M MajorTOM 23TB model outperforms PhilEO Globe 0.5TB 44M. We also show that for most n-shots for road density estimation and building density regression, PhilEO 200M FastTOM outperforms all the other models. The effectiveness of both dataset and model scaling is validated using the PhilEO Bench. We also study the impact of architecture scaling, transitioning from U-Net Convolutional Neural Networks (CNN) to Vision Transformers (ViT).
@@
Earth Observation, Foundation Models.
@@ I Introduction
Given the vast amounts of satellite data coming from multiple constellations, such as Copernicus Sentinel-2, the focus has recently been shifted to Self-Supervised Learning (SSL). Masked Autoencoders [1] and Contrastive Learning [2] aim at reducing the need for large labeled datasets and have led to a growing interest in Foundation Models (FM) for Earth Observation (EO) [4], [5]. FMs are pretrained on satellite data to learn robust feature representations and, then, are fine-tuned on smaller labeled datasets for various downstream tasks. FMs use different architectures for the encoding (and decoding) of the data and features, which lead to different performance results for the downstream tasks, e.g. pixel-wise regression.
In this work, we focus not only on scaling the volume of the pretraining data, but also on scaling the number of parameters and the architecture, studying the transition from CNN-based models to Vision Transformer (ViT) backbones combined with UPerNet decoder. The main contributions of this paper are:
@@
  • We scale the PhilEO pretraining from the PhilEO Globe 0.5TB dataset to the full MajorTOM 23TB dataset, as well as to the specialized FastTOM 2TB subset (i.e. without oceans and ice), with different number of model parameters and architectures [^1] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We demonstrate that larger pretraining data reduces the Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$ for road density regression for the $n$-shot $n=500$.
@@
  • In addition to dataset scaling, we study the scaling of the model parameters and the architecture, comparing CNN- and ViT-based models. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering road and building density estimation and land cover mapping.
@@
  • We develop and compare different architectures for the EO Foundation Models in order to select which one to scale-up. For road density regression, we show that the PhilEO ViT UPerNet outperforms the PhilEO ViT with a convolutional decoder in RMSE, with a percentage improvement of $34.2\%$ for the $n$-shot $n=100$. On the same task, the PhilEO Geo-Aware U-Net model achieves the same performance as the PhilEO ViT UPerNet.
@@
The rest of the paper is organized as follows. Section II provides background on the PhilEO Geospatial Foundation Model (GFM) and the PhilEO Bench evaluation, as well as the datasets MajorTOM and FastTOM. Section III describes the proposed PhilEO MajorTOM 23TB and FastTOM 2TB models. Section IV presents the related work. We evaluate our models on the PhilEO Bench in Sec. V for road and building density estimation and land cover mapping. We present the results of the PhilEO MajorTOM 23TB and FastTOM 2TB models, the PhilEO 200M FastTOM, the PhilEO Geo-Aware U-Net, and the PhilEO ViT UPerNet 100M FastTOM.
@@ II Background
Figure 1: High-level overview of ViT-UPerNet. The dimensions of the feature map $C_{4}$ are maintained. The other feature maps are upsampled, respectively downsampled. The resulting feature map is used for downstream processing.
PhilEO Bench. To enable fair evaluation and comparison of GFMs, [3] introduced the PhilEO Bench which includes a global labeled Sentinel-2 dataset for the standardized downstream tasks: road density estimation, building density regression, and land cover mapping. The bench also incorporates several existing GFMs, such as [6] and [7], allowing for consistent comparisons across models.
MajorTOM and FastTOM datasets. The MajorTOM dataset [15] comprises approximately 60TB of unlabeled global data, covering most of the Earth, including oceans and ice. In this work, we focus on the MajorTOM Core-S2L2A subset, 23TB. FastTOM is a 2TB specialized subset of MajorTOM Core-S2L2A, containing only land and excluding oceans and ice, making it more task-specific for terrestrial downstream tasks.
@@ III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
We extend our previously proposed PhilEO model by pretraining it on the MajorTOM Core-S2L2A dataset [15], which contains 23TB data covering the vast majority of the Earth’s surface, including land, oceans, and ice. To the best of the authors’ knowledge, this is the first instance where the full 23TB MajorTOM dataset is used to pretrain a GFM. The PhilEO MajorTOM model was trained using the Leonardo Davinci-1 Supercomputer, using between four and eight compute nodes, each equipped with four NVIDIA A100 GPUs (40GB VRAM), resulting in a total of 16-32 GPUs. This HPC configuration provided approximately 2.5–5 petaFLOPS of compute performance. Without such infrastructure, training on a single GPU would have taken several months of continuous computation, highlighting the critical role of HPC resources in enabling this work. To handle the scale of the dataset and model, we used PyTorch’s Distributed Data Parallel (DDP) to accelerate training. For larger ViT-based models still ongoing training on MajorTOM, we are transitioning to Fully Sharded Data Parallel (FSDP) to improve memory efficiency and support larger model scales. In addition, mixed-precision training (BF16/BF32) was used for improved throughput and reduced memory usage. Alongside the MajorTOM pretraining, we also trained a model variant on the FastTOM 2TB subset, which contains only land. FastTOM serves as a proxy for MajorTOM: strong performance on FastTOM is expected to translate into good performance when scaling to the full MajorTOM 23TB dataset. This strategy allows for efficient experimentation and architecture selection before undertaking full-scale pretraining.
@@ III-B PhilEO ViT UPerNet
PhilEO Bench [3] standardizes model comparison by using a common convolutional decoder based on a U-Net-like architecture [8]. In this work, we introduce an alternative decoder strategy by implementing the UPerNet decoder [10], shown in Fig. 1, in the PhilEO Bench, in order to compare models with a ViT backbone. The UPerNet design [10] draws inspiration from human visual perception, combining hierarchical feature extraction through a Feature Pyramid Network (FPN) [21] with global context aggregation via a Pyramid Pooling Module (PPM). The FPN aggregates features across multiple scales through a top-down pathway with lateral connections, while the PPM captures context at various spatial resolutions. ViT-UPerNet architectures have been shown to achieve state-of-the-art performance in various computer vision tasks [11]-[14], including segmentation and classification. For models using a ViT backbone, we extract intermediate feature maps, $\{C_{2},C_{3},C_{4},C_{5}\}$ in Fig. 1, across different layers. These feature maps are then appropriately resized (upsampled or downsampled) to align spatial dimensions. The highest-level feature map $C_{5}$ is processed by the PPM, which applies pooling operations at multiple scales to capture contextual information at different resolutions. The outputs of the PPM and the FPN are then fused through a series of 1×1 convolutions and element-wise additions, producing the final set of multi-scale feature maps, $\{P_{2},P_{3},P_{4},P_{5}\}$ in Fig. 1. These aggregated features are then used for downstream pixel-wise prediction tasks, including regression and segmentation.

### v2

# Earth Observation Foundation Model PhilEO: Pretraining on the MajorTOM and FastTOM Datasets

@@ Abstract
Today, Earth Observation (EO) satellites generate massive volumes of data. To fully exploit this, it is essential to pretrain Foundation Models (FMs) on large unlabeled datasets, enabling efficient fine-tuning for downstream tasks with minimal labeled data. We study scaling-up FMs, where scaling refers to incorporating a larger pretraining dataset that includes both oceans and ice in addition to land (MajorTOM, 23TB), a larger model in terms of number of parameters (200M), and an architecture that achieves top performance, efficiency, and low computational complexity, being capable of handling images covering large areas. We develop various models using different architectures, including U-Net Convolutional Neural Networks (CNN), Vision Transformers (ViT), and Mamba State-Space Models (SSM), and different numbers of parameters. We evaluate the FLoating-point OPerations (FLOPs) needed by the models. ViT, due to attention, suffers from quadratic complexity, while Mamba achieves linear complexity, scaling-up better. We present the scaling of the model PhilEO on MajorTOM, covering the vast majority of Earth, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. We fine-tune the models on the PhilEO Bench for different downstream tasks: roads, buildings, and land cover. For all n-shots for roads, the Geo-Aware U-Net 44M-23T model outperforms U-Net 44M-0.5T. For most n-shots for roads and buildings, U-Net 200M-2T outperforms the other models. We show that using Mamba models, we achieve comparable results on the downstream tasks, with less computational expenses. We also compare with the recent FM TerraMind which we evaluate on the PhilEO Bench.
@@
Earth Observation, Foundation Models, remote sensing, Vision Transformer (ViT), Mamba state-space model.
@@ I Introduction
Given the vast amounts of satellite data coming from multiple constellations, such as Copernicus Sentinel-2, the focus has recently been shifted to Self-Supervised Learning (SSL). Today, Earth Observation (EO) satellites generate massive volumes of data, with the Copernicus Sentinel-2 constellation alone producing approximately 1.6TB per day. To fully exploit this information captured by remote sensing satellites, it is essential to pretrain EO Foundation Models (FMs) using SSL on large unlabeled datasets, enabling efficient fine-tuning for various downstream tasks with minimal labeled data.
In this work, our main aim is to study the scaling-up of EO FMs, where the scaling-up refers to incorporating a larger pretraining dataset that includes both oceans and ice in addition to land, for example the MajorTOM dataset 23TB (Sentinel-2 L2A, Core-S2L2A), a larger model in terms of number of parameters, for example 200M, and an architecture that is suitable for scaling-up in terms of performance, efficiency, speed, and computational complexity, being capable of handling high-resolution data and images that cover large areas. Regarding the latter, i.e. regarding the architecture we select to scale-up, we note that the Vision Transformer (ViT), because of attention, suffers from quadratic complexity with respect to the image token sequence length. On the contrary, the recent Mamba S6 State-Space Model (SSM) achieves linear complexity and scales more favorably to high-resolution data and images that cover large geographical areas [30]. In this way, global correlations and long-range dependencies within the EO image that covers a large area can be effectively and efficiently modelled. Moreover, U-Net Convolutional Neural Network (CNN) models also have linear complexity, but cannot model global correlations and long-range dependencies. In addition, in this work, scaling-up also refers to including the time aspect in the pretraining, for example seasonality and the four seasons of summer, winter, autumn, and fall, as well as additional modalities, for example Sentinel-1 (S-1) SAR data in addition to Sentinel-2 (S-2) multi-spectral images.
@@
SSL in EO and remote sensing. Masking Auto-Encoder (MAE) [1] and contrastive learning [2] to pretrain EO FMs using SSL aim at reducing the need for large labeled datasets and have led to a growing interest in FMs for EO [4], [5]. EO FMs are pretrained on satellite data to learn robust feature representations and, then, are fine-tuned on smaller labeled datasets for various downstream tasks. EO FMs use different architectures for the encoding (and the decoding) of the data and features, which leads to different performance results for the downstream tasks, e.g. for pixel-wise regression [3].
EO FMs can be trained, for example, using MAE [1], [56], [57], contrastive learning [2], or Joint Embedding Predictive Architecture (JEPA) [39]. MAE refers to masking (e.g., a large percentage of the image such as $75\%$), using the reconstruction loss, and the MSE [23], [26]. Also, the ViT [9] (or a variant of the ViT like the Swin Transformer [12]) is usually used. Here, we also note that the Spectral Angle Mapper (SAM), which takes into account the spectral signatures of the observations, can also be used in addition to the MSE, or even on its own as the loss function that is minimized during training. In addition, contrastive learning refers to performing data augmentation and bringing closer together in the latent feature space the data-augmented samples, repelling the dissimilar samples.
Furthermore, JEPA [40] (or I-JEPA [41]) refers to performing reconstruction in the latent feature space, estimating and predicting in the feature space the representation of a part of the image based on another part of the image (e.g., similar to masking), and operating in the latent space rather than the pixel image space. The feature space is closer to encoding semantics and is more robust than pixels because, for example, in S-2, the image space may have clouds or missing data.
Our scaling-up study. In this paper, we present the scaling-up of our recently proposed model PhilEO on the unlabeled 23TB dataset MajorTOM [15], which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. FastTOM is a task-specific dataset for land only. In the literature, there exist FMs only for oceans or ice, for example HydroFM [25], but we are interested in Earth Observation and not only in Land Observation or Ocean Observation (or Ice Observation). To this end, in this work, we develop and study various models with different numbers of parameters, and also with different architectures. We evaluate and examine the FLoating-point OPerations (FLOPs) that are needed by the models. We fine-tune the models on the PhilEO Bench [3] for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance.
Models of size 44M and 200M, and different architectures. In this paper, in the beginning, we first develop and examine models using different architectures pretrained on the PhilEO Globe 0.5TB dataset. Next, we scale-up the best of the examined architectures by increasing the number of parameters and pretraining the models with bigger datasets, i.e. MajorTOM and FastTOM. We evaluate the performance of these models on the examined downstream tasks, and we find the best models. Also, we compare the models with the RS3Mamba model [48] which comprises a Mamba encoder, a CNN auxiliary encoder, and a custom multi-scale decoder. Our results demonstrate that for most $n$-shots for road density estimation and building density regression, Geo-Aware U-Net 200M-2TB outperforms all the other models we examine. We also show that for all $n$-shots for the task of road density regression, the PhilEO Geo-Aware U-Net 44M-23TB model outperforms Geo-Aware U-Net 44M-0.5TB. The effectiveness of both dataset and model scaling of the different EO FMs is validated using the PhilEO Bench [3]. We also study the impact of architecture scaling, transitioning from U-Net CNN to ViT, where the UPerNet decoder brings significant performance gains, as well as to 2D Mamba models [31].
Jointly increasing the pretraining dataset size and the model size. In this paper, we focus not only on scaling-up the volume of the pretraining data (i.e. to 23TB) [15], but also on scaling the number of parameters and the architecture, also studying the transition from CNN-based U-Net models to ViT backbones combined with the UPerNet decoder, as well as to Mamba for images. Here, the recent Mamba S6 SSM is a potential effective paradigm-shift solution [30], as it combines global dependencies modeling and linear computational scaling, and this is important. The self-attention mechanism of the ViT is not sensitive to the token order, while on the contrary, the selectivity mechanism of Mamba (i.e. the selective scan) is sensitive to the token order. Notably, it needs a large hidden state dimension as it has an exponentially decaying recall accuracy based on the distance from the current token [36]. In addition, ViT has quadratic complexity with respect to the image token sequence length, due to the attention Query, Key, Value mechanism, which is based on the scaled dot product similarity measure, while the 2D Mamba S6 SSM model is faster and more efficient as it achieves linear complexity with respect to the image token sequence length [30], [31].
In light of all the above, this paper’s main contributions are:
@@
  • We scale the pretraining from 0.5TB to 23TB, which is a big scale-up, as well as to 2TB, where the 23TB are general-purpose EO data (i.e. MajorTOM [^1] http://github.com/ESA-PhiLab/Major-TOM and also in http://huggingface.co/Major-TOM), which include land, oceans, and ice, while the 2TB are a specialized subset that does not include oceans and ice. We use different number of model parameters and architectures [^2] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We show that for the task of road density regression, for the $n$-shot $n=500$, larger pretraining data reduces the final downstream task’s Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$.
@@
  • In addition to pretraining dataset scaling, we also study the scaling-up of the model parameters and the architecture, comparing CNN- and ViT-based models, as well as Mamba. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering both road and building density estimation, as well as land cover mapping.
@@
  • We develop and compare different architectures for the EO Foundation Models in order to select which one to scale-up. More specifically, for road density regression, we show that the ViT UPerNet model outperforms the ViT with a convolutional decoder in RMSE, with a percentage improvement of $34.2\%$ for the $n$-shot $n=100$. On the same task, the U-Net CNN Geo-Aware model achieves the same performance as the ViT UPerNet.
Further contributions of this paper. In addition, in this work, because scaling-up also refers to including additional EO modalities, for example S-1 SAR in addition to S-2, we release and make publicly available the new S12-PhilEOBench dataset for fine-tuning and downstream tasks. S12-PhilEOBench is for evaluating EO Foundation Models that are multi-modal and use co-located S-2 and S-1 data.
@@
Organization of this paper. The rest of the paper is organized as follows. Section II provides background on the PhilEO Geospatial Foundation Model (GFM) and the PhilEO Bench evaluation, as well as the datasets MajorTOM 23TB and FastTOM 2TB. Section III describes the proposed model. Section IV presents the related work. We evaluate the models on the PhilEO Bench evaluation framework in Sec. V for road and building density estimation, and land cover mapping. We present the results of the PhilEO MajorTOM 23TB and FastTOM 2TB models, as well as the results of the PhilEO Geo-Aware U-Net, the PhilEO ViT UPerNet 100M MajorTOM and FastTOM, and the PhilEO 200M FastTOM.
@@ II Background
PhilEO Bench. To enable fair evaluation and comparison of GFMs, [3] introduced the PhilEO Bench, which includes a global labeled Sentinel-2 dataset for the standardized downstream tasks: road density estimation, building density regression, and land cover mapping. The bench also incorporates several existing GFMs, such as [6] and [7], allowing for consistent comparisons across models.
Figure 1: High-level overview of ViT UPerNet. The dimensions of the feature map $C_{4}$ are maintained. The other feature maps are upsampled, respectively downsampled. The resulting feature map is used for downstream processing.
MajorTOM and FastTOM datasets. The MajorTOM dataset [15] comprises in total approximately 60TB of unlabeled global data, covering most of the Earth, including oceans and ice. In this work, we focus on the MajorTOM Core-S2L2A subset, 23TB. In addition, FastTOM is a 2TB specialized subset of MajorTOM Core-S2L2A, containing only land and excluding oceans and ice, making it more task-specific for terrestrial downstream tasks.
@@ III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
We extend our previously proposed PhilEO model by pretraining it on the MajorTOM Core-S2L2A dataset [15], which contains 23TB data covering the vast majority of the Earth’s surface, including land, oceans, and ice. To the best of the authors’ knowledge, this is the first instance where the full 23TB MajorTOM dataset is used to pretrain a GFM.
The PhilEO MajorTOM model was trained using the Leonardo Davinci-1 Supercomputer, using between four and eight compute nodes, each equipped with four NVIDIA A100 GPUs (40GB VRAM), resulting in a total of 16-32 GPUs. This HPC configuration provided approximately 2.5–5 petaFLOPS of compute performance, which is important. Without such infrastructure, training on a single GPU would have taken several months of continuous computation, thus highlighting the critical role of HPC resources in enabling this work.
To handle the scale of the dataset and model, we used PyTorch’s Distributed Data Parallel (DDP) to accelerate training. For larger ViT-based models still ongoing training on MajorTOM, we are transitioning to Fully Sharded Data Parallel (FSDP) to improve memory efficiency and support even larger model scales. In addition, mixed-precision training (BF16/BF32) was used for the training, for improved throughput and reduced memory usage.
Alongside the MajorTOM pretraining, we also trained a model variant on the FastTOM 2TB subset, which contains only land. Here, we note that FastTOM serves as a proxy for MajorTOM: strong performance on FastTOM is expected to translate into good performance when scaling-up to the full MajorTOM 23TB dataset. This strategy allows for efficient experimentation and architecture selection before undertaking the expensive full-scale pretraining.
@@ III-B PhilEO ViT UPerNet
PhilEO Bench [3] standardizes model comparison by using a common convolutional decoder based on a U-Net-like architecture [8]. In this work, we introduce an alternative decoder strategy by implementing the UPerNet decoder [10], shown in Fig. 1, in the PhilEO Bench, in order to compare models with a ViT backbone. The UPerNet design [10] draws inspiration from human visual perception, combining hierarchical feature extraction through a Feature Pyramid Network (FPN) [21] with global context aggregation via a Pyramid Pooling Module (PPM). The FPN aggregates features across multiple scales through a top-down pathway with lateral connections, while the PPM captures context at various spatial resolutions.
ViT-UPerNet architectures have been shown to achieve state-of-the-art performance in various computer vision tasks [11]-[14], including segmentation and classification. For models that use a ViT backbone, we extract intermediate feature maps, $\{C_{2},C_{3},C_{4},C_{5}\}$ in Fig. 1, across different layers. These feature maps are then appropriately resized (i.e. upsampled or downsampled) to align spatial dimensions.
The highest-level feature map $C_{5}$ is processed by the PPM, which applies pooling operations at multiple scales to capture contextual information at different resolutions. The outputs of the PPM and the FPN are then fused through a series of 1×1 convolutions and element-wise additions, producing the final set of multi-scale feature maps, $\{P_{2},P_{3},P_{4},P_{5}\}$ in Fig. 1. These aggregated features are then used for downstream pixel-wise prediction tasks, including regression and segmentation.
III-C Mamba 2D Models
In this paper, our main aim is to study scaling-up and we include in our analysis Mamba S6 2D SSM models [36]. We develop, train, and evaluate models based on Mamba and we include this type of models (i.e. Mamba) in the PhilEO Bench, and this is important. Mamba models are able to handle long sequences and have demonstrated strong performance in tasks that require long-range correlations and dependencies [38]. We note that this type of models has not been tested before in the evaluation framework PhilEO Bench, and also not in other evaluation testbeds such as Geo-Bench, Copernicus-Bench, FoMo-Bench [49], PANGAEA, EarthNets, and SustainBench.
We develop, train, and evaluate Mamba models that use input-dependent matrices (i.e. the S6 model compared to, for example, the S4 SSM) [33]. The Mamba models we train and test are for images, i.e. 2D, so that there are limited spatial discrepancies introduced due to the Mamba model being sensitive to the image token order. More specifically, we do not use 1D Mamba, which is for 1D causal signals, because this would introduce spatial discrepancies between neighbouring pixels in an image based on the scan traversal strategy that is used (i.e. based on how the image is unrolled), and this is important. To use 1D Mamba models in images, a reverse scan can be used [30], [55], that is, because we focus on spatial dependencies and correlation (and not, for example, on temporal correlation), the Mamba models for images, i.e. 2D, are bidirectional. Notably, regarding an example of not a 2D Mamba model, 1D Mamba would not have spatial continuity (and also not as strong correlations and dependencies) for the top and left adjacent pixels (i.e. and also for the entire $8$-pixel neighbourhood of each pixel) in an EO image [37], [36].
Figure 2: Example images from S12-PhilEOBench from East Africa where we have, from the left, S-2, co-located S-1 SAR GRD $V$, and S-1 GRD VH.
Figure 3: Further example images from S12-PhilEOBench where we have, starting from the left, S-2 multi-spectral and co-located S-1 SAR GRD $V$ and VH.
For the Mamba 2D models which we will also examine in Sec. V, they are based on continuous-time dynamics [30], [33]. We denote the state by $h$, and for the linear dynamics (i.e. similar to the Kalman filter), we have:
$h^{\prime}(t)=\textbf{A}h(t)+\textbf{B}x(t)$ (1)
$y(t)=\textbf{C}h(t)+\textbf{Z}x(t)$ (2)
where the input signal is denoted by $x$ and the output by $y$. The matrices A, C, B, and Z are the state transition matrix, the output matrix, the input matrix, and the input affecting the output matrix, respectively [54]. Here, we note that the input affecting the output matrix, Z, is the residual connections in the network. The state, $h$, is of dimension $N$, i.e. $h\in\mathbb{R}^{N}$. In addition, $x\in\mathbb{R}$ and $y\in\mathbb{R}$. Moreover, for the matrices, we also have: $\textbf{A}\in\mathbb{R}^{N\times N}$, $\textbf{C}\in\mathbb{R}^{1\times N}$, $\textbf{B}\in\mathbb{R}^{N\times 1}$, and $\textbf{Z}\in\mathbb{R}$.
The next steps are discretization [31], [30], making the matrices input-dependent (i.e. S6, Mamba), and also ensuring spatial continuity for the 2D images (i.e. for EO data) rather than introducing spatial discrepancies due to the traversal and unrolling of the image to a vector [38]. For the discretization of the continuous-time dynamics in (1) and (2), we have:
$h_{k}=\bar{\textbf{A}}h_{k-1}+\bar{\textbf{B}}x_{k}$ (3)
$y_{k}=\textbf{C}h_{k}$ (4)
where because of using residual connections, we omit the input affecting the output matrix. The main advantage of the Mamba S6 SSM model is the linear complexity with respect to the input sequence length (in contrast, for example, to the ViT which has quadratic complexity), and the models which we will also examine in Sec. V are for images, i.e. 2D [36]-[38].
Mamba models, instead of attention and the Query, Key, Value multi-head mechanism [9], use selectivity. For the linear dynamics in (3) and (4), the input-dependent matrices are:
$\bar{\textbf{A}}\,=\,\exp(\textbf{S}_{\Delta}\textbf{A})$ (5)
$\bar{\textbf{B}}\,=\,(\textbf{S}_{\Delta}\textbf{A})^{-1}\,(\exp(\textbf{S}_{\Delta}\textbf{A})-\textbf{I})\,\textbf{S}_{\Delta}\textbf{S}_{\textbf{B}}$ (6)
where $\textbf{S}_{\textbf{B}}$ is the output of the learnable linear projection for making the matrix B input-dependent, i.e. $\textbf{S}_{\textbf{B}}=\textbf{x}\textbf{W}_{\textbf{B}}^{T}$, where $\textbf{x}\in\mathbb{R}^{B\times L\times D}$ is the input, $B$ is the batch size, $L$ is the image token sequence length, and $D$ is the embedding dimension [30], [36]. In addition, the learnable linear projection is denoted by $\textbf{W}_{\textbf{B}}\in\mathbb{R}^{N\times D}$. Here, $\textbf{S}_{\textbf{B}}\in\mathbb{R}^{B\times L\times N}$. Similarly, for making the matrix C input-dependent, we have: $\textbf{C}=\textbf{x}\textbf{W}_{\textbf{C}}^{T}$, where we denote the learnable linear projection by $\textbf{W}_{\textbf{C}}\in\mathbb{R}^{N\times D}$. Also, $\textbf{C}\in\mathbb{R}^{B\times L\times N}$ [30], [31]. We note that for ViT, attention is also based on learnable linear projections for the Query, Key, Value matrices that are input-dependent.
Furthermore, in (5) and (6), $\Delta$ is the sampling interval of the continuous parameters and $\textbf{S}_{\Delta}\in\mathbb{R}^{B\times L\times D}$. $\Delta$ is an interval, i.e. the step size for the continuous dynamics [33], [51]. We note that using Mamba S6 SSM models, we are able to use, leverage, train, and fit the learnable model parameter $\Delta$ which controls the sampling rate for the continuous parameters, and this is important. The parameter $\Delta$ of Mamba [52], [53] is like a forgetting mechanism that is able to control how much a specific sample is forgotten or remembered. In particular, if $\Delta\to\infty$, then the model focuses more on the input, while if $\Delta\to 0$, the focus is more on the previous state. For the input-dependent $\Delta$ and its learnable linear projection, we have:
$\textbf{S}_{\Delta}\,=\,f(\textbf{x}\textbf{W}_{\Delta}^{T})\text{,}$ (7)
$\text{where the function $f(.)$ is: }f(x)=\tau_{\Delta}\,\text{Broadcast}_{D}(x)$ (8)
where $\textbf{W}_{\Delta}\in\mathbb{R}^{1\times D}$ is the learnable linear projection (i.e., fully-connected (FC) layer) for making the time step $\Delta$ input-dependent [30], [63]. In addition, $\textbf{S}_{\Delta}$ is a function of the output of the learnable linear projection, $\textbf{x}\textbf{W}_{\Delta}^{T}$. Moreover, in (7) and (8), $\tau_{\Delta}$ is the softplus function. The $\text{Broadcast}_{D}(.)$ operation broadcasts the result to all the $D$ dimensions, i.e. from $1$ to $D$, as $\textbf{S}_{\Delta}\in\mathbb{R}^{B\times L\times D}$.